Geometry. The lenght of a rectangle is 2 in. more than twice its width. If the perimeter of the rectangel is 34 in., find the dimensions of the rectangle.
Answer by rapaljer(4671)

Let x = width
2x+2 = length

The perimeter is 34 inches.

2W + 2L = 34
2(___)+2(____)=34
2(x) + 2(2x+2) = 34
2x + 4x + 4 = 34
6x + 4 = 34
6x+4-4 = 34 - 4
6x = 30
x=5 Width
2x+2 = Length
2(5)+2 = 12 Length

Check:
2W + 2L = P
2(5)+2(12) = 34
10 +24 = 34

By the way, I have a TON of these problems in a Lesson Plan in algebra.com. See also my own website by clicking on my tutor name "rapaljer" anywhere in algebra.com. Then go to the "MATH IN LIVING COLOR" page, and look for "Basic Algebra", then "Section 1.10 Word Problems"!!

R^2 Retired from SCC